3) You are given a solution of 14.8 M NH3. How many milliliter of this solution uire to give 100.0 mL of 1.00 M NH3 when diluted?
Answer #1

Number of moles of NH3 = molarity of NH3 * volume of solution in L

Number of moles of NH3 = 1.00 * 0.1 L = 0.1 mole

Molarity = number of moles / volume of solution in L

14.8 = 0.1 mole / volume of solution in L

volume of solution in L = 0.1 / 14.8 = 0.00676 L

1 L = 1000 mL

0.00676 L = 6.76 mL

Therefore, the volume of solution is 6.76 mL